### v1

# A Weighted Llarull Type Theorem

Abstract.
This paper generalizes Llarull’s classical scalar curvature rigidity theorem to the setting of weighted manifolds with P-scalar curvature. More precisely, We obtain the P-scalar curvature rigidity of the round sphere. Furthermore, we prove the improvement of Llarull’s theorem due to Listing [15] for P-scalar curvature.
Key words and phrases:
Degree, Weighted Dirac Operator, Index Theorem, Weighted Scalar Curvature.
2020 Mathematics Subject Classification:
53C21, 53C24, 53C27
1. Introduction
The famous scalar curvature rigidity theorem of the round sphere was proved by Llarull in [16], which confirms the Gromov’s conjecture in [7]. For a comprehensive overview of extremality and rigidity properties of Riemannian manifolds, we refer to Gromov’s Four lectures on scalar curvature [8]. In fact, Llarull established the following theorem.
Theorem 1.1.
([16, Theorem C]) Let $(M,g,)$ be a closed spin Riemannian manifold of dimension $n\geq 3$ with $R(g)\geq n(n-1)$. If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth, area non-increasing map of non-zero degree, then $h$ is an isometry.
This theorem reveals the interrelationship among metric, scalar curvature, and topology. As an improvement of Llarull’s theorem, Listing generalized previous results by omitting the assumption of $\|\wedge^{2}h_{*}\|\leq 1$ and using the function $\|\wedge^{2}h_{*}\|$ in the scalar curvature inequality as follows.
Theorem 1.2.
([15, Theorem 1 ]) Let $(M^{n},g)$ be a closed spin Riemannian manifold of dimension $n\geq 3$. If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth map of non-zero degree such that $R(g)\geq n(n-1)\|\wedge^{2}h_{*}\|$, then there exists a constant $c>0$ such that $h:(M,c\cdot g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is an isometry.
Here $\|\wedge^{2}h_{*}\|$ denotes the operator norm of the second exterior power of the differential of $h$, see Definition 1.3. For the proof of this theorem, we also refer to Theorem 3.1 in [20, Theorem 3.1].
The proofs of Theorem 1.1 and Theorem 1.2 all rely on the Dirac operator method, requiring the hypothesis that M is spin. An open question is whether the spin assumption can be removed in both Theorem 1.1 and Theorem 1.2. Recently, Cecchini-Wang-Xie-Zhu [4] proved Theorem 1.1 without the spin assumption in 4-dimension provided $h$ is a distance non-increasing map.
A weighted Riemannian manifold is a Riemannian manifold $(M,g)$ equipped with a smooth function $f:M\to R$ and a measure $e^{-f}dV_{g}$ on $M$. It was first introduced by Lichnerowicz in [13, 14]. Weighted manifolds have appeared for a long time in differential geometry. They play an important role in Perelman’s approach to the Ricci flow. A crucial idea of Perelman’s spectacular proofs [17] required considering manifolds with density and their evolution. This led him to introduce the notion of weighted scalar curvature
$R_{f}(g)\coloneqq R(g)+2\Delta_{g}f-|\nabla f|^{2}_{g}\,$,
which is not the trace of the weighted Ricci curvature of Bakry-Émery, sometimes called the P-scalar curvature. Weighted manifolds with positive or non-negative Weighted scalar curvature have very similar properties compared to Riemannian manifolds with positive or non-negative scalar curvature. By the Gauss-Bonnet theorem and the divergence theorem, every oriented closed surface with positive Weighted scalar curvature must be a topological sphere. In recent years, there have been numerous studies on weighted scalar curvature. Schoen-Yau [31] showed that if $(M^{3},g)$ is oriented with positive scalar curvature, then it contains no closed, immersed stable minimal surfaces of positive genus, Fan [6] generalized this result to the weighted case (see also[3, 9] for subsequent work). Baldauf-Ozuch [2] proved the positive weighted mass theorem under the spin assumption by generalizing Witten’s spinor argument in [21] to the weighted case. Chu-Zhu [5] proved the positive weighted mass theorem under the dimension $3\leq n\leq 7$ assumption by generalizing Schoen-Yau’s minimal surface method in [18, 19] to the weighted case. For more works of the weighted scalar curvature supporting the above philosophy, we refer the reader to [1, 10, 11, 12].
To state our main theorem, we need to introduce some definitions.
Definition 1.3.
Suppose $h:(M,g)\to(N,\bar{g})$ is a smooth map between Riemannian manifolds $(M,g)$ and $(N,\bar{g})$.
  • (i)
    The operator norm of the differential of $h$ is defined as
    $\|h_{*}\|:M\to[0,+\infty),\,p\,\to\max_{0\neq v\in T_{p}M}\left[\frac{h^{*}\bar{g}(v,v)}{g(v,v)}\right]^{\frac{1}{2}}.$
  • (ii)
    The operator norm of the second exterior power of the differential of $h$ is defined as
    $\|\wedge^{2}h_{*}\|:M\to[0,+\infty),\,p\,\to\max_{0\neq\eta\in\wedge^{2}T_{p}M}\left[\frac{h^{*}\bar{g}(\eta,\eta)}{g(\eta,\eta)}\right]^{\frac{1}{2}}.$
Then $h$ is said to be distance non-increasing if $\|h_{*}\|\leq 1$ and $h$ is said to be area non-increasing if $\|\wedge^{2}h_{*}\|\leq 1$.
Remark 1.4.
The area scaling can be bounded by the square of the distance scaling:
$0\leq\|\wedge^{2}h_{*}\|\leq\|h_{*}\|^{2}.$
Our main results generalize Theorem 1.1 and Theorem 1.2 to the weighted case.
Theorem 1.5.
Let $(M^{n},g,e^{-f}\mathrm{d}V_{g})$ be a weighted closed spin Riemannian manifold of dimension $n\geq 3$. If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth map of non-zero degree such that $R_{f}\geq n(n-1)\|\wedge^{2}h_{*}\|$, then there exists a constant $c>0$ such that $h:(M,c\cdot g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is an isometry. Consequently, $f$ is a constant.
Remark 1.6.
The proof of Theorem 1.5 follows Llarull’s approach and using the Poincaré type inequality in [20]. The key step of our proof is to solve the weighted function $f$ and the operator norm $\|\wedge^{2}h_{*}\|$ by constructing a perfect square expression and using integration by parts in a differential equation.
As a corollary, we have the following weighted Llarull type theorem.
Theorem 1.7.
Let $(M,g,e^{-f}\mathrm{d}V_{g})$ be a weighted closed spin Riemannian manifold of dimension $n\geq 3$ with $R_{f}\geq n(n-1)$. If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth, area non-increasing map of non-zero degree, then $h$ is an isometry. Consequently, $f$ is a constant.
Remark 1.8.
Excluding Theorem 1.5, one can follow Llarull’s approach by using the weighted Dirac operator method to prove Theorem 1.7 without essential difficulties.
The paper is organized as follows. In section 2, we give some facts about the spin manifolds and weighted Dirac operator. In section 3, we prove Theorem 1.5.
Acknowledgements: The second author would like to thank Professor Yukai Sun for much help and many discussions.
2. Spin manifold and Weighted Dirac Operator
In this section, we recall some key facts concerning spin manifolds and weighted Dirac operator.
A spin manifold $M^{n}$ is an oriented manifold with a spin structure on its tangent bundle $TM$. Note that a manifold $M$ is spin if and only if the first and second Stiefel-Whitney class of tangent bundle vanish, namely $w_{1}(TM)=w_{2}(TM)=0$. There is a complex spinor bundle $S\coloneqq S(TM)$ with its Hermitian metric $\langle,\rangle_{S}$ and compatible connection $\nabla^{S}$ over a spin Riemannian manifold $(M,g)$. The Dirac operator is a first order elliptic differential operator acting on spinors and is given by
$D\coloneqq\sum_{i=1}^{n}e_{i}\cdot\nabla^{S}_{e_{i}}\,$,
where $\{e_{i}\}_{i=1}^{n}$ is a local orthonormal frame of $TM$ and $\cdot$ denotes the Clifford multiplication. For a twisted spinor bundle $S\otimes E$, the twisted Dirac operator $D_{E}:\Gamma(S\otimes E)\to\Gamma(S\otimes E)$ is defined as
$D_{E}\coloneqq\sum_{i=1}^{n}e_{i}\cdot\nabla^{S\otimes E}_{e_{i}}\,$,
where $E$ is a given vector bundle over $M$.
Let $(M,g)$ be a spin Riemannian manifold and $f\in C^{\infty}(M)$. The weighted Dirac operator $D_{f,E}:\Gamma(S\otimes E)\to\Gamma(S\otimes E)$ is defined as
$D_{f,E}\coloneqq D_{E}-\frac{1}{2}\nabla f\cdot,$
where $D_{E}$ is the standard twisted Dirac operator defined above, $\cdot$ denotes the Clifford multiplication. For instance, let $\varphi\otimes v\in\Gamma(S\otimes E)$, then
$D_{f,E}(\varphi\otimes v)=D_{E}(\varphi\otimes v)-\frac{1}{2}(\nabla f\cdot\varphi)\otimes v\,.$
As standard Dirac operator, the weighted Dirac operator is self-adjoint with respect to the weighted $L^{2}$-inner product. By slight abuse of notation, we denote the connection on $S\otimes E$ by $\nabla$.
Proposition 2.1.
The weighted Dirac operator $D_{f,E}$ is self-adjoint, i.e.,
$\int_{M}\langle D_{f,E}\Psi,\Phi\rangle e^{-f}=\int_{M}\langle\Psi,D_{f,E}\Phi\rangle e^{-f},$
where $\Psi,\Phi\in\Gamma(S\otimes E)$ and $\langle,\rangle$ is the induce metic on $S\otimes E$ .
Proof.
We fix an arbitrary point $p\in M$ and work with a local orthonormal frame $\{e_{i}\}$ such that $\nabla e_{i}=0$ at $p$. Then, at $p$, for any $\Psi,\Phi\in\Gamma(S\otimes E),$
$\langle D_{f,E}\Psi,\Phi\rangle$ $=\langle D_{E}\Psi-\frac{1}{2}\nabla f\cdot\Psi,\Phi\rangle$
$=\langle e_{i}\cdot\nabla_{e_{i}}\Psi,\Phi\rangle-\frac{1}{2}\langle\nabla f\cdot\Psi,\Phi\rangle$
$=-\langle\nabla_{e_{i}}\Psi,e_{i}\cdot\Phi\rangle+\frac{1}{2}\langle\Psi,\nabla f\cdot\Phi\rangle$
$=-e_{i}\langle\Psi,e_{i}\cdot\Phi\rangle+\langle\Psi,e_{i}\cdot\nabla_{e_{i}}\Phi\rangle+\frac{1}{2}\langle\Psi,\nabla f\cdot\Phi\rangle$
$=\langle\Psi,e_{i}\cdot\nabla_{e_{i}}\Phi\rangle-\frac{1}{2}\langle\Psi,\nabla f\cdot\Phi\rangle-e_{i}\langle\Psi,e_{i}\cdot\Phi\rangle+\langle\Psi,\nabla f\cdot\Phi\rangle$
$=\langle\Psi,D_{f,E}\Phi\rangle-e_{i}\langle\Psi,e_{i}\cdot\Phi\rangle+\langle\Psi,\nabla f\cdot\Phi\rangle.$
The last two terms are a weighted divergence term. In fact, define a vector field $X\in\Gamma(TM\otimes\mathbb{C})$ on $M$ by
$g(X,Y)\coloneqq-\langle\Psi,Y\cdot\Phi\rangle$
for all vector field $Y\in\Gamma(TM)$. Then (again computing at a fixed arbitrary point p)
$\mathrm{div}_{f}X$ $\coloneqq\mathrm{div}X-g(X,\nabla f)$
$=g(\nabla_{e_{i}}X,e_{i})-g(X,\nabla f)$
$=e_{i}g(X,e_{i})-g(X,\nabla f)$
$=-e_{i}\langle\Psi,e_{i}\cdot\Phi\rangle+\langle\Psi,\nabla f\cdot\Phi\rangle.$
The proposition follows from integration by parts. ∎
Similarly, the weighted Laplacian $\Delta_{f}:\Gamma(S\otimes E)\to\Gamma(S\otimes E)$ is defined as
$\Delta_{f}\coloneqq-\nabla^{*}\nabla-\nabla_{\nabla f},$
it is also self-adjoint with respect to the weighted $L^{2}$-inner product.
Proposition 2.2.
The weighted Laplacian $\Delta_{f}$ is a nonpositive operator and is self-adjoint. In fact,
$\int_{M}\langle\Delta_{f}\Psi,\Phi\rangle e^{-f}=-\int_{M}\langle\nabla\Psi,\nabla\Phi\rangle e^{-f}=\int_{M}\langle\Psi,\Delta_{f}\Phi\rangle e^{-f}.$
Proof.
Once again, we fix an arbitrary point $p\in M$ and work with a local orthonormal frame $\{e_{i}\}$ such that $\nabla e_{i}=0$ at $p$. Then, at $p$,
$\langle\Delta_{f}\Psi,\Phi\rangle$ $=\langle\nabla_{e_{i}}\nabla_{e_{i}}\Psi,\Phi\rangle-\langle\nabla_{\nabla_{f}}\Psi,\Phi\rangle$
$=-\langle\nabla_{e_{i}}\Psi,\nabla_{e_{i}}\Phi\rangle+e_{i}\langle\nabla_{e_{i}}\Psi,\Phi\rangle-\langle\nabla_{\nabla_{f}}\Psi,\Phi\rangle$
$=-\langle\nabla\Psi,\nabla\Phi\rangle+e_{i}\langle\nabla_{e_{i}}\Psi,\Phi\rangle-\langle\nabla_{\nabla_{f}}\Psi,\Phi\rangle.$
As before, the last two terms are a weighted divergence term. In fact, define a vector field $U\in\Gamma(TM\otimes\mathbb{C})$ on $M$ by
$g(U,V)\coloneqq\langle\nabla_{V}\Psi,\Phi\rangle$
for all vector field $V\in\Gamma(TM)$. Then
$\mathrm{div}_{f}U=e_{i}\langle\nabla_{e_{i}}\Psi,\Phi\rangle-\langle\nabla_{\nabla_{f}}\Psi,\Phi\rangle.$
The desired formula follows by integration by parts. ∎
Similar to the standard Lichnerowicz formula, a weighted Lichnerowicz formula holds, which was first observed by Perelman [17]. Now, we state and prove it here.
Lemma 2.3 (weighted Lichnerowicz formula).
Let $(M^{n},g,e^{-f}\mathrm{d}V_{g})$ be a weighted closed spin Riemannian manifold, then
$D_{f,E}^{2}=-\Delta_{f}+\frac{1}{4}R_{f}+\mathcal{R}^{E},$
where $\mathcal{R}^{E}$ is defined on simple elements $\sigma\otimes v\in\Gamma(S\otimes E)$ by $\mathcal{R}^{E}(\sigma\otimes v)=\frac{1}{2}\Sigma_{i,j=1}^{n}\left(e_{i}\cdot e_{j}\cdot\sigma\right)\otimes R^{E}(e_{i},e_{j})v$, $R^{E}$ is the curvature tensor of $E$.
Proof.
For any $\Phi\in\Gamma(S\otimes E)$, we have
$D^{2}_{f,E}\Phi=$ $\left(D_{E}-\frac{1}{2}\nabla f\cdot\right)\left(D_{E}-\frac{1}{2}\nabla f\cdot\right)\Phi$
$=$ $D_{E}^{2}\Phi-\frac{1}{2}\nabla f\cdot\left(e_{i}\cdot\nabla_{e_{i}}\Phi\right)-\frac{1}{2}\left(e_{i}\cdot\nabla_{e_{i}}\right)\left(\nabla f\cdot\Phi\right)-\frac{1}{4}\|\nabla f\|^{2}\Phi$
$=$ $D_{E}^{2}\Phi-\frac{1}{2}\nabla f\cdot e_{i}\cdot\nabla_{e_{i}}\Phi-\frac{1}{2}e_{i}\cdot\langle\nabla_{e_{i}}\nabla f,e_{j}\rangle e_{j}\cdot\Phi$
$-\frac{1}{2}e_{i}\cdot\nabla f\cdot\nabla_{e_{i}}\Phi-\frac{1}{4}\|\nabla f\|^{2}\Phi$
$=$ $D_{E}^{2}\Phi-\frac{1}{2}\left(e_{i}\cdot\nabla f+\nabla f\cdot e_{i}\right)\cdot\nabla_{e_{i}}\Phi-\frac{1}{2}\mathrm{Hess}f(e_{i},e_{j})e_{i}\cdot e_{j}\cdot\Phi$
$-\frac{1}{4}\|\nabla f\|^{2}\Phi$
$=$ $\left(\nabla^{*}\nabla+\frac{1}{4}R+\mathcal{R}^{E}\right)\Phi+\langle\nabla f,e_{i}\rangle\nabla_{e_{i}}\Phi+\frac{1}{2}\left(\Delta f\right)\Phi-\frac{1}{4}\|\nabla f\|^{2}\Phi$
$=$ $\left(\nabla^{*}\nabla+\nabla_{\nabla f}\right)\Phi+\frac{1}{4}\left(R+2\Delta f-\|\nabla f\|^{2}\right)\Phi+\mathcal{R}^{E}\Phi$
$=$ $\left(-\Delta_{f}+\frac{1}{4}R_{f}+\mathcal{R}^{E}\right)\Phi.$
We have used the standard Lichnerowicz formula and the properties of Clifford multiplication. ∎
3. proof of Theorem 1.5
In this section, We adopt Llarull’s setting to prove the main theorem in two cases: even-dimensional case and odd-dimensional case as described in [16].
Proof of Theorem1.5.
Let $\{e_{i}\}_{i=1}^{n}$ be a $g$-orthonormal frame near $p\in M$ and $\{\epsilon_{i}\}_{i=1}^{n}$ be a $g_{\mathbb{S}^{n}}$-orthonormal frame near $h(p)\in\mathbb{S}^{n}$. Moreover, we can require that there exist non-negative scalars $\{\lambda_{i}\}_{i=1}^{n}$ such that $h_{*}(e_{i})=\lambda_{i}\epsilon_{i}$. We will show that
(3.1) $R_{f}$ $=n(n-1)\|\wedge^{2}h_{*}\|$
and
(3.2) $\lambda_{i}$ $=\|\wedge^{2}h_{*}\|^{\frac{1}{2}}$
for $1\leq i\leq n$.
[…]
$\Phi=a_{\alpha\beta}\sigma_{\alpha}\otimes v_{\beta}.$
By Proposition 2.2 and weighted Lichnerowicz formula, we have
$\int_{M}\langle D_{f,E}^{2}\Phi,\Phi\rangle e^{-f}=$ $\int_{M}\langle\nabla\Phi,\nabla\Phi\rangle e^{-f}+\frac{1}{4}\int_{M}R_{f}\|\Phi\|^{2}e^{-f}+\int_{M}\langle\mathcal{R}^{E}\Phi,\Phi\rangle e^{-f}.$
According to Lemma 4.5 in [16], the curvature term has the following estimate,
$\langle\mathcal{R}^{E}\Phi,\Phi\rangle\geq-\frac{1}{4}\sum_{i\neq j=1}^{2m}\lambda_{i}\lambda_{j}\|\Phi\|^{2}.$
Consequently,
$\int_{M}\langle D_{f,E}^{2}\Phi,\Phi\rangle e^{-f}\geq\int_{M}\|\nabla\Phi\|^{2}e^{-f}+\frac{1}{4}\int_{M}(R_{f}-\sum_{i\neq j=1}^{2m}\lambda_{i}\lambda_{j})\|\Phi\|^{2}e^{-f}.$
Next, we prove $R_{f}=2m(2m-1)\|\wedge^{2}h_{*}\|$ by contradiction. If $R_{f}(x)>2m(2m-1)\|\wedge^{2}h_{*}\|(x)$ for some $x\in M$, then $ker(D_{f,E}^{2})=0$, by Proposition 2.1, $ker(D_{f,E})=0$. The weighted Dirac operator $D_{f,E}:\Gamma((S\otimes E^{+})\oplus(S\otimes E^{-}))\to\Gamma((S\otimes E^{+})\oplus(S\otimes E^{-}))$ preserves the direct sum, so $D_{f,E^{+}}\coloneqq D_{f,E}|_{S\otimes E^{+}}$ has zero kernel, i.e. $ker(D_{f,E^{+}})=0$. We also have that $D_{f,E^{+}}=D^{+}_{f,E^{+}}\oplus D^{-}_{f,E^{+}}$, where $D^{\pm}_{f,E^{+}}:\Gamma(S^{\pm}\otimes E^{+})\to\Gamma(S^{\mp}\otimes E^{+})$. Since $0=ker(D_{f,E^{+}})=ker(D^{+}_{f,E^{+}})\oplus ker(D^{-}_{f,E^{+}})$, we have $ker(D^{+}_{f,E^{+}})=ker(D^{-}_{f,E^{+}})=0$. Therefore the index of $D^{+}_{f,E^{+}}$ is given by
$Index(D^{+}_{f,E^{+}})=ker(D^{+}_{f,E^{+}})-ker(D^{-}_{f,E^{+}})=0.$
We can define an elliptic operator $D^{+}_{f,E^{+}}(t)\coloneqq D^{+}_{E^{+}}-\frac{t}{2}\nabla f\cdot$ for $0\leq t\leq 1$. It is easy to see that the weighted Dirac operator $D^{+}_{f,E^{+}}$ is homotopic to the standard Dirac operator $D^{+}_{E^{+}}$, where $D^{+}_{E^{+}}:\Gamma(S^{\pm}\otimes E^{+})\to\Gamma(S^{\mp}\otimes E^{+})$. Since the index is homotopy invariant, we can deduce that $Index(D^{+}_{E^{+}})=0$. But Atiyah-Singer index theorem states that $Index(D^{+}_{E^{+}})\neq 0$ (it is equals to Euler characteristic of $\mathbb{S}^{2m}$ times degree of $h$), which leads to a contradiction. Therefore, $R_{f}\equiv 2m(2m-1)\|\wedge^{2}h_{*}\|$ on $M$. As $Index(D^{+}_{f,E^{+}})\neq 0,\,ker(D_{f,E})\neq 0$, there exists $0\neq\Psi\in\Gamma(S\otimes E)$ such that $D_{f,E}\Psi=0$ and
$0\geq\int_{M}\|\nabla\Psi\|^{2}e^{-f}+\frac{1}{4}\int_{M}\left(2m(2m-1)\|\wedge^{2}h_{*}\|-\sum_{i\neq j=1}^{2m}\lambda_{i}\lambda_{j}\right)\|\Psi\|^{2}e^{-f},$
combine with $\lambda_{i}\lambda_{j}\leq\|\wedge^{2}h_{*}\|$ for $1\leq i\neq j\leq 2m$, we have $\lambda_{i}\lambda_{j}=\|\wedge^{2}h_{*}\|$ for $1\leq i\neq j\leq 2m$. Therefore, $\lambda_{i}=\|\wedge^{2}h_{*}\|^{\frac{1}{2}}$ for $1\leq i\leq 2m$.
Odd Dimensional Case $n=2m-1\geq 3$: Let $M\times\mathbb{S}_{r}^{1}$ be the Riemannian product of $M$ with the circle $\mathbb{S}_{r}^{1}$ of radius $r\geq 1$. Consider the following map
$(M\times\mathbb{S}_{r}^{1},g+r^{2}ds^{2})\xrightarrow{h\times\frac{\mathrm{id}}{r}}(\mathbb{S}^{2m-1}\times\mathbb{S}^{1},g_{\mathbb{S}^{2m-1}}+ds^{2})\xrightarrow{\,\gamma\,}\mathbb{S}^{2m-1}\wedge\mathbb{S}^{1}\cong\mathbb{S}^{2m},$
where $h\times\frac{\mathrm{id}}{r}$ is given by $(h\times\frac{\mathrm{id}}{r})(x,t)=(h(x),\frac{t}{r})$ for all $(x,t)\in M\times\mathbb{S}_{r}^{1}$, and $\gamma$ is a distance non-increasing map of nonzero degree.
Define $\tilde{h}=\gamma\circ(h\times\frac{\mathrm{id}}{r})$. Consider two complex spinor bundles $S$ over $M\times\mathbb{S}_{r}^{1}$ and $E_{0}$ over $\mathbb{S}^{2m}$, respectively. And consider the bundle $S\otimes E$ over $M\times\mathbb{S}_{r}^{1}$ for $E=\tilde{h}^{*}E_{0}$.
Let $\{\tilde{e}_{i}\}_{i=1}^{2m}$ be a $g+r^{2}ds^{2}$-orthonormal frame near $(p,\theta)\in M\times\mathbb{S}_{r}^{1}$ such that $\tilde{e}_{1},...,\tilde{e}_{2m-1}$ are tangent to $M$ and $\tilde{e}_{2m}$ is tangent to $\mathbb{S}_{r}^{1}$. Let $\{\tilde{\epsilon}_{i}\}_{i=1}^{2m}$ be a $g_{\mathbb{S}^{2m}}$-orthonormal frame near $\tilde{h}(p,\theta)\in\mathbb{S}^{2m}$. Moreover, we can require that there exist non-negative scalars $\{\tilde{\lambda}_{i}\}_{i=1}^{2m}$ such that $\tilde{h}_{*}(\tilde{e}_{i})=\tilde{\lambda}_{i}\tilde{\epsilon}_{i}$. Note that
$\tilde{\lambda}_{i}^{2}=g_{\mathbb{S}^{2m}}(\tilde{h}_{*}(\tilde{e}_{i}),\tilde{h}_{*}(\tilde{e}_{i}))\leq g_{\mathbb{S}^{2m-1}}(h_{*}(\tilde{e}_{i}),h_{*}(\tilde{e}_{i}))\leq\|h_{*}\|^{2}\leq K^{2},$
$\tilde{\lambda}_{i}^{2}\tilde{\lambda}_{j}^{2}=g_{\mathbb{S}^{2m}}(\tilde{h}_{*}(\tilde{e}_{i}\wedge\tilde{e}_{j}),\tilde{h}_{*}(\tilde{e}_{i}\wedge\tilde{e}_{j}))\leq g_{\mathbb{S}^{2m-1}}(h_{*}(\tilde{e}_{i}\wedge\tilde{e}_{j}),h_{*}(\tilde{e}_{i}\wedge\tilde{e}_{j}))\leq\|\wedge^{2}h_{*}\|^{2}$
and
$\tilde{\lambda}_{2m}^{2}=g_{\mathbb{S}^{2m}}(\tilde{h}_{*}(\tilde{e}_{2m}),\tilde{h}_{*}(\tilde{e}_{2m}))\leq ds^{2}(\frac{\tilde{e}_{2m}}{r},\frac{\tilde{e}_{2m}}{r})=\frac{1}{r^{4}}\leq\frac{1}{r^{2}},$
where $K$ is a positive constant independent of $r$ and $1\leq i\neq j\leq 2m-1$. Thus, $0\leq\tilde{\lambda}_{i}\leq\|h_{*}\|\leq K$, $0\leq\tilde{\lambda}_{i}\tilde{\lambda}_{j}\leq\|\wedge^{2}h_{*}\|$ and $0\leq\tilde{\lambda}_{2m}\leq\frac{1}{r}$.
As before, according to Lemma 4.5 in [16], we have
$\begin{split}\langle\mathcal{R}^{E}\Phi,\Phi\rangle=&\frac{1}{4}\sum_{i\neq j=1}^{2m-1}\sum_{k,l}\sum_{\alpha,\beta}\tilde{\lambda}_{i}\tilde{\lambda}_{j}a_{\alpha\beta}a_{kl}\langle e_{i}\cdot e_{j}\cdot\sigma_{\alpha}\otimes\epsilon_{j}\cdot\epsilon_{i}\cdot v_{\beta},\sigma_{k}\otimes v_{l}\rangle\\ &+\frac{1}{4}\sum_{i=1}^{2m-1}\sum_{k,l}\sum_{\alpha,\beta}\tilde{\lambda}_{i}\tilde{\lambda}_{2m}a_{\alpha\beta}a_{kl}\langle e_{i}\cdot e_{2m}\cdot\sigma_{\alpha}\otimes\epsilon_{2m}\cdot\epsilon_{i}\cdot v_{\beta},\sigma_{k}\otimes v_{l}\rangle\\ &+\frac{1}{4}\sum_{j=1}^{2m-1}\sum_{k,l}\sum_{\alpha,\beta}\tilde{\lambda}_{2m}\tilde{\lambda}_{j}a_{\alpha\beta}a_{kl}\langle e_{2m}\cdot e_{j}\cdot\sigma_{\alpha}\otimes\epsilon_{j}\cdot\epsilon_{2m}\cdot v_{\beta},\sigma_{k}\otimes v_{l}\rangle\\ &\geq-\frac{1}{4}\sum_{i\neq j=1}^{2m-1}\tilde{\lambda}_{i}\tilde{\lambda}_{j}\|\Phi\|^{2}-\frac{1}{2}\sum_{i=1}^{2m-1}\tilde{\lambda}_{i}\tilde{\lambda}_{2m}\|\Phi\|^{2}\\ &\geq-\frac{1}{4}\sum_{i\neq j=1}^{2m-1}\tilde{\lambda}_{i}\tilde{\lambda}_{j}\|\Phi\|^{2}-\frac{(2m-1)K}{2r}\cdot\|\Phi\|^{2}.\end{split}$
Consequently,
(3.3) $\begin{split}\int_{M\times\mathbb{S}_{r}^{1}}\langle D_{f,E}^{2}\Phi,\Phi\rangle e^{-f}\geq&\int_{M\times\mathbb{S}_{r}^{1}}\|\nabla\Phi\|^{2}e^{-f}+\frac{1}{4}\int_{M\times\mathbb{S}_{r}^{1}}R_{f}\|\Phi\|^{2}e^{-f}\\ &-\frac{1}{4}\int_{M\times\mathbb{S}_{r}^{1}}\left(\sum_{i\neq j=1}^{2m-1}\tilde{\lambda}_{i}\tilde{\lambda}_{j}+\frac{2(2m-1)K}{r}\right)\|\Phi\|^{2}e^{-f}.\end{split}$
Next, we follow [20] to prove $R_{f}=(2m-1)(2m-2)\|\wedge^{2}h_{*}\|$ by contradiction. Assume to the contrary that the inequality $R_{f}\geq(2m-1)(2m-2)\|\wedge^{2}h_{*}\|$ is strict somewhere. Then there exists $x_{0}\in M$ and two positive constants $\delta,\delta_{0}>0$ such that for any $y\in U_{\delta}(x_{0})$ with
$\left(R_{f}-(2m-1)(2m-2)\|\wedge^{2}h_{*}\|\right)(y)>\delta_{0},$
where $U_{\delta}(x_{0})$ is a $\delta$-neighborhood of $x_{0}$ in $M$.
[…]
(3.4) $\int_{M\times\mathbb{S}_{r}^{1}}\|\Psi\|^{2}e^{-f}\leq C\int_{U_{\delta}(x_{0})\times\mathbb{S}_{r}^{1}}\|\Psi\|^{2}e^{-f}+C\int_{M\times\mathbb{S}_{r}^{1}}\|\nabla\Psi\|^{2}e^{-f},$
here we take the metric $e^{-f}\cdot\langle\,$$\rangle$ on spinor bundle $S\otimes E$ and $C$ is a positive constant which is independent of $r$.
Since $D_{f,E}\Psi=0$, combining equation (3.3) with equation (3.4), it follows that
$\begin{split}0\geq&\int_{M\times\mathbb{S}_{r}^{1}}\|\nabla\Psi\|^{2}e^{-f}+\frac{1}{4}\int_{M\times\mathbb{S}_{r}^{1}}\left(R_{f}-\sum_{i\neq j=1}^{2m-1}\tilde{\lambda}_{i}\tilde{\lambda}_{j}\right)\|\Psi\|^{2}e^{-f}\\ &-\frac{2m-1}{2r}\int_{M\times\mathbb{S}_{r}^{1}}\|{h}_{*}\|\|\Psi\|^{2}e^{-f}\\ \geq&\int_{M\times\mathbb{S}_{r}^{1}}\|\nabla\Psi\|^{2}e^{-f}+\frac{\delta_{0}}{4}\int_{U_{\delta}(x_{0})\times\mathbb{S}_{r}^{1}}\|\Psi\|^{2}e^{-f}\\ &-\frac{(2m-1)K}{2r}\int_{M\times\mathbb{S}_{r}^{1}}\|\Psi\|^{2}e^{-f}\\ \geq&\left(1-\frac{(2m-1)KC}{2r}\right)\int_{M\times\mathbb{S}_{r}^{1}}\|\nabla\Psi\|^{2}e^{-f}\\ &+\left(\frac{\delta_{0}}{4}-\frac{(2m-1)KC}{2r}\right)\int_{U_{\delta}(x_{0})\times\mathbb{S}_{r}^{1}}\|\Psi\|^{2}e^{-f}.\end{split}$
Since $C$ and $K$ are independent of $r$, for a given sufficiently large $r$, the above estimates imply that $\Psi$ vanishes on $U_{\delta}(x_{0})\times\mathbb{S}_{r}^{1}$ and $\nabla\Psi$ vanishes on $M\times\mathbb{S}_{r}^{1}$. This together with the inequality (3.4) implies that $\Psi$ vanishes on $M\times\mathbb{S}_{r}^{1}$, which leads to a contradiction. Hence, we have proved that $R_{f}=(2m-1)(2m-2)\|\wedge^{2}h_{*}\|$. By the same argument, we also have
$\sum_{i\neq j=1}^{2m-1}\tilde{\lambda}_{i}\tilde{\lambda}_{j}=(2m-1)(2m-2)\|\wedge^{2}h_{*}\|,$
combining with $\tilde{\lambda}_{i}\tilde{\lambda}_{j}\leq\|\wedge^{2}h_{*}\|$ for $1\leq i\neq j\leq 2m-1$, we see that $\tilde{\lambda}_{i}\tilde{\lambda}_{j}=\|\wedge^{2}h_{*}\|$ for $1\leq i\neq j\leq 2m-1$. Suppose for $i<j$ that $\tilde{e}_{i}\wedge\tilde{e}_{j}=\sum_{\bar{i}<\bar{j}}T_{ij}^{\bar{i}\bar{j}}e_{\bar{i}}\wedge e_{\bar{j}}$ near $p\in M$, note that $\sum_{\bar{i}<\bar{j}}\left(T_{ij}^{\bar{i}\bar{j}}\right)^{2}=1$ since $\tilde{e}_{i}\wedge\tilde{e}_{j}$ is a unit 2-vector. We have
$\|\wedge^{2}h_{*}\|^{2}=\tilde{\lambda}_{i}^{2}\tilde{\lambda}_{j}^{2}$ $=g_{\mathbb{S}^{2m}}(\tilde{h}_{*}(\tilde{e}_{i}\wedge\tilde{e}_{j}),\tilde{h}_{*}(\tilde{e}_{i}\wedge\tilde{e}_{j}))$
$\leq g_{\mathbb{S}^{2m-1}}(h_{*}(\tilde{e}_{i}\wedge\tilde{e}_{j}),h_{*}(\tilde{e}_{i}\wedge\tilde{e}_{j}))$
$=\sum_{\bar{i}<\bar{j}}\sum_{\bar{k}<\bar{l}}T_{ij}^{\bar{i}\bar{j}}T_{ij}^{\bar{k}\bar{l}}g_{\mathbb{S}^{2m-1}}(h_{*}(e_{\bar{i}}\wedge e_{\bar{j}}),h_{*}(\tilde{e}_{\bar{k}}\wedge\tilde{e}_{\bar{l}}))$
$=\sum_{\bar{i}<\bar{j}}\sum_{\bar{k}<\bar{l}}T_{ij}^{\bar{i}\bar{j}}T_{ij}^{\bar{k}\bar{l}}\lambda_{\bar{i}}\lambda_{\bar{j}}\lambda_{\bar{k}}\lambda_{\bar{l}}\delta_{\bar{i}\bar{k}}\delta_{\bar{j}\bar{l}}$
$=\sum_{\bar{i}<\bar{j}}\left(T_{ij}^{\bar{i}\bar{j}}\right)^{2}\lambda_{\bar{i}}^{2}\lambda_{\bar{j}}^{2}$
$\leq\|\wedge^{2}h_{*}\|^{2}.$
Thus, $\lambda_{i}\lambda_{j}=\|\wedge^{2}h_{*}\|$ for $1\leq i\neq j\leq 2m-1$, which implies that $\lambda_{i}=\|\wedge^{2}h_{*}\|^{\frac{1}{2}}$ for $1\leq i\leq 2m-1$.
So far, we have proved that $R_{f}=n(n-1)\|\wedge^{2}h_{*}\|$ and $\lambda_{i}=\|\wedge^{2}h_{*}\|^{\frac{1}{2}}$ for $1\leq i\leq n$. Define $u\coloneqq\|\wedge^{2}h_{*}\|=\lambda_{i}^{2}$ and $U=\{x\in M:u(x)\neq 0\}$, note that $u$ is a smooth function on $M$. Since $h$ has non-zero degree, $U$ is non-empty. Thus, $h^{*}g_{\mathbb{S}^{n}}=u\cdot g$ on $U$. By the formula of scalar curvature under conformal change, we have on $U$
(3.5) $n(n-1)=h^{*}[R(g_{\mathbb{S}^{n}})]=u^{-1}R(g)-(n-1)u^{-2}\Delta u-\frac{(n-1)(n-6)}{4}u^{-3}|\nabla u|^{2}.$
Since $R_{f}(g)=R(g)+2\Delta f-|\nabla f|^{2}=n(n-1)u$, $u$ and $\nabla u$ vanish on $M\setminus U$, we can deduce that
(3.6) $u^{k}e^{lf}\left(-2u^{2}\Delta f+u^{2}|\nabla f|^{2}-(n-1)u\Delta u-\frac{1}{4}(n-1)(n-6)|\nabla u|^{2}\right)=0$
on the entire $M$ for all $k\geq 0$ and $l\in R$. By integration by parts, we have
(3.7) $\begin{split}0=&(1+2l)\int_{M}u^{k+2}e^{lf}|\nabla f|^{2}\\ &+(n-1)\left(k+1-\frac{n-6}{4}\right)\int_{M}u^{k}e^{lf}|\nabla u|^{2}\\ &+(2k+4+ln-l)\int_{M}u^{k+1}e^{lf}g(\nabla u,\nabla f).\end{split}$
By taking
(3.8) $l=\frac{-n^{2}+(11+4\sqrt{2})n-10-4\sqrt{2}}{2(n-1)^{2}}~~~\text{and}~~k=\frac{n-2}{4}\geq 0,$
we have
$1+2l=$ $\left(\frac{1+2\sqrt{2}}{\sqrt{n-1}}\right)^{2},$
$2k+4+ln-l=$ $2(4+\sqrt{2}),$
$(n-1)\left(k+1-\frac{n-6}{4}\right)=$ $2(n-1).$
It follows from equation (3.7) taht
$\int_{M}u^{\frac{n-2}{4}}e^{lf}|Au\nabla f+B\nabla u|^{2}=0,$
where $A=\frac{1+2\sqrt{2}}{\sqrt{n-1}}$, $B=\sqrt{2(n-1)}$. Thus, $u^{\frac{n-2}{4}}|Au\nabla f+B\nabla u|^{2}=0$ on $M$. Moreover, we have
(3.9) $Au\nabla f+B\nabla u\equiv 0~~~\text{on}\,\,M.$
Consequently,
[…]
Plug $u\nabla f=-\frac{B}{A}\nabla u$ and equation (3.10) into equation (3.6) for k=0,m=0k=0,\,m=0, we obtain
$\left(2\frac{B}{A}-(n-1)\right)u\Delta u+\left(\frac{B^{2}}{A^{2}}-2\frac{B}{A}-\frac{1}{4}(n-1)(n-6)\right)|\nabla u|^{2}=0,$
by integration by parts, we see that
$\frac{1}{196}(n-1)\left((23-32\sqrt{2})n-30+144\sqrt{2}\right)\int_{M}|\nabla u|^{2}=0.$
Therefore, $\nabla u\equiv 0.$ Since $h$ has non-zero degree, $u$ is a positive constant on M, i.e. $u\equiv c>0.$ It follows that $h:(M,c\cdot g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a local isometry. Since $M$ is closed and $\mathbb{S}^{n}$ is simply connected, we conclude that $h:(M,c\cdot g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is an isometry. $f$ is a constant comes from equation (3.9). This completes the proof. ∎
Remark 3.1.
For $n\geq 2$, the Theorem 1.5 also holds if one replaces the operator norm $\|\wedge^{2}h_{*}\|$ by $\|h_{*}\|^{2}$.

### v2

@@
-A Weighted Llarull Type Theorem
+A Weighted Llarull Type Theorem and its applications
 Abstract.
 This paper generalizes Llarull’s classical scalar curvature rigidity theorem to the setting of weighted manifolds with P-scalar curvature.
-More precisely, We obtain the P-scalar curvature rigidity of the round sphere.
-Furthermore, we prove the improvement of Llarull’s theorem due to Listing [15] for P-scalar curvature.
+More precisely, we prove the refinement of Llarull’s theorem for P-scalar curvature, which is similar to Listing’s work [20].
+As an application, we establish a Llarull type theorem in the form of $\mathbb{S}^{k}\times\mathbb{T}^{n-k}$.
 Key words and phrases:
@@
 Theorem 1.1.
-([16, Theorem C]) Let $(M,g,)$ be a closed spin Riemannian manifold of dimension $n\geq 3$ with $R(g)\geq n(n-1)$.
+([21, Theorem C]) Let $(M,g,)$ be a closed spin Riemannian manifold of dimension $n\geq 3$ with $R(g)\geq n(n-1)$.
 If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth, area non-increasing map of non-zero degree, then $h$ is an isometry.
@@
 Theorem 1.2.
-([15, Theorem 1 ]) Let $(M^{n},g)$ be a closed spin Riemannian manifold of dimension $n\geq 3$.
+([20, Theorem 1 ]) Let $(M^{n},g)$ be a closed spin Riemannian manifold of dimension $n\geq 3$.
 If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth map of non-zero degree such that $R(g)\geq n(n-1)\|\wedge^{2}h_{*}\|$, then there exists a constant $c>0$ such that $h:(M,c\cdot g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is an isometry.
 Here $\|\wedge^{2}h_{*}\|$ denotes the operator norm of the second exterior power of the differential of $h$, see Definition 1.3.
-For the proof of this theorem, we also refer to Theorem 3.1 in [20, Theorem 3.1].
+For the proof of this theorem, we also refer to Theorem 3.1 in [25, Theorem 3.1].
 The proofs of Theorem 1.1 and Theorem 1.2 all rely on the Dirac operator method, requiring the hypothesis that M is spin.
 An open question is whether the spin assumption can be removed in both Theorem 1.1 and Theorem 1.2.
-Recently, Cecchini-Wang-Xie-Zhu [4] proved Theorem 1.1 without the spin assumption in 4-dimension provided $h$ is a distance non-increasing map.
-A weighted Riemannian manifold is a Riemannian manifold $(M,g)$ equipped with a smooth function $f:M\to R$ and a measure $e^{-f}dV_{g}$ on $M$.
+Recently, Cecchini-Wang-Xie-Zhu [6] proved Theorem 1.1 without the spin assumption in 4-dimension provided $h$ is a distance non-increasing map, Theorem 1.2 plays an important role in their proof.
+A weighted Riemannian manifold is a Riemannian manifold $(M,g)$ equipped with a smooth function $f:M\to\mathbb{R}$ and a measure $e^{-f}dV_{g}$ on $M$.
 It was first introduced by Lichnerowicz in [18, 19].
@@
 This led him to introduce the notion of weighted scalar curvature
-$R_{f}(g)\coloneqq R(g)+2\Delta_{g}f-|\nabla f|^{2}_{g}\,$,
+$R_{f}(g)\coloneqq R(g)+2\Delta_{g}f-|\nabla_{g}f|^{2}_{g}\,$,
 which is not the trace of the weighted Ricci curvature of Bakry-Émery, sometimes called the P-scalar curvature.
@@
 The key step of our proof is to solve the weighted function $f$ and the operator norm $\|\wedge^{2}h_{*}\|$ by constructing a perfect square expression and using integration by parts in a differential equation.
-As a corollary, we have the following weighted Llarull type theorem.
-Theorem 1.7.
-Let $(M,g,e^{-f}\mathrm{d}V_{g})$ be a weighted closed spin Riemannian manifold of dimension $n\geq 3$ with $R_{f}\geq n(n-1)$.
+Corollary 1.7.
+([9, Theorem 1.3]) Let $(M,g,e^{-f}\mathrm{d}V_{g})$ be a weighted closed spin Riemannian manifold of dimension $n\geq 3$ with $R_{f}\geq n(n-1)$.
 If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth, area non-increasing map of non-zero degree, then $h$ is an isometry.
@@
 Remark 1.8.
-Excluding Theorem 1.5, one can follow Llarull’s approach by using the weighted Dirac operator method to prove Theorem 1.7 without essential difficulties.
-The paper is organized as follows. In section 2, we give some facts about the spin manifolds and weighted Dirac operator. In section 3, we prove Theorem 1.5.
+Deng has already proved Corollary 1.7 for general weighted scalar curvature in [9].
+As a corollary of Theorem 1.5, we obtain the Llarull type theorem also holds for warped $\mathbb{T}^{k}$-extension of $M$.
+see also results of Bär-Brendle-Hanke-Wang in [3] and Wang-Xie in [25] for the case with a non-empty boundary.
+Theorem 1.9.
+Let $(M^{n},g)$ be a spin Riemannian manifold of dimension $n\geq 3$, the warped $\mathbb{T}^{k}$-extension of $M$ is the manifold $(M^{n}\times\mathbb{T}^{k},\bar{g})$ with $\bar{g}=g+\sum_{i=1}^{k}e^{-2f_{i}}d\theta^{2}$, where $f_{i}$ is a smooth function on $M$ for each $1\leq i\leq k$.
+Assume that $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth map of non-zero degree such that $R(\bar{g})\geq n(n-1)\|\wedge^{2}h_{*}\|$, then there exists a constant $c>0$ such that $h:(M,c\cdot g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is an isometry.
+Consequently, $f_{i}$ is a constant for each $1\leq i\leq k$.
+Remark 1.10.
+For $n\geq 2$, Theorem 1.5 and Theorem 1.9 also hold if one replaces the operator norm $\|\wedge^{2}h_{*}\|$ by $\|h_{*}\|^{2}$.
+As an application of Theorem 1.9, we generalize Theorem 1.1 in [13] to $\mathbb{S}^{k}$ for $4\leq k\leq 6$ by following the approach of Hao-Shi-Sun in [13].
+Theorem 1.11.
+Suppose $4\leq n\leq 7$ and $3\leq k\leq n-1$.
+Let $(M^{n},g)$ be a closed spin Riemannian manifold.
+If $h:(M,g)\to(\mathbb{S}^{k}\times\mathbb{T}^{n-k},g_{\mathbb{S}^{k}}+g_{\mathbb{T}^{n-k}})$ is a smooth map of non-zero degree such that $R(g)\geq k(k-1)\|\wedge^{2}h_{*}\|$, then there exists a constant $c>0$ such that $(M,g)$ is locally isometric to $(\mathbb{S}^{k}\times\mathbb{T}^{n-k},\frac{1}{c}\cdot(g_{\mathbb{S}^{k}}+g_{\mathbb{T}^{n-k}}))$.
+Here, we would like to make a statement: The Theorem 1.11 in the case of $h$ is area non-increasing was also obtained independently by Chow in a recent preprint [7].
+The paper is organized as follows.
+In section 2, we give some facts about the spin manifolds and weighted Dirac operator.
+In section 3, we prove Theorem 1.5.
+In section 4, we prove Theorem 1.9 and Theorem 1.11.
 Acknowledgements: The second author would like to thank Professor Yukai Sun for much help and many discussions.
@@
 $D_{f,E}(\varphi\otimes v)=D_{E}(\varphi\otimes v)-\frac{1}{2}(\nabla f\cdot\varphi)\otimes v\,.$
-As standard Dirac operator, the weighted Dirac operator is self-adjoint with respect to the weighted $L^{2}$-inner product.
+As standard Dirac operator, the weighted Dirac operator is also self-adjoint with respect to the weighted $L^{2}$-inner product.
 By slight abuse of notation, we denote the connection on $S\otimes E$ by $\nabla$.
@@
 Proposition 2.2.
-The weighted Laplacian $\Delta_{f}$ is a nonpositive operator and is self-adjoint. In fact,
+The weighted Laplacian $\Delta_{f}$ is a non-positive operator and is self-adjoint. In fact,
 $\int_{M}\langle\Delta_{f}\Psi,\Phi\rangle e^{-f}=-\int_{M}\langle\nabla\Psi,\nabla\Phi\rangle e^{-f}=\int_{M}\langle\Psi,\Delta_{f}\Phi\rangle e^{-f}.$
@@
 3. Proof of Theorem 1.5
-In this section, We adopt Llarull’s setting to prove the main theorem in two cases: even-dimensional case and odd-dimensional case as described in [16].
+In this section, We adopt Llarull’s setting to prove Theorem 1.5 in two cases: even-dimensional case and odd-dimensional case as described in [21].
 Proof of Theorem1.5.
@@
 (3.3) $\begin{split}\int_{M\times\mathbb{S}_{r}^{1}}\langle D_{f,E}^{2}\Phi,\Phi\rangle e^{-f}\geq&\int_{M\times\mathbb{S}_{r}^{1}}\|\nabla\Phi\|^{2}e^{-f}+\frac{1}{4}\int_{M\times\mathbb{S}_{r}^{1}}R_{f}\|\Phi\|^{2}e^{-f}\\ &-\frac{1}{4}\int_{M\times\mathbb{S}_{r}^{1}}\left(\sum_{i\neq j=1}^{2m-1}\tilde{\lambda}_{i}\tilde{\lambda}_{j}+\frac{2(2m-1)K}{r}\right)\|\Phi\|^{2}e^{-f}.\end{split}$
-Next, we follow [20] to prove $R_{f}=(2m-1)(2m-2)\|\wedge^{2}h_{*}\|$ by contradiction.
+Next, we follow Wang-Xie [25] to prove $R_{f}=(2m-1)(2m-2)\|\wedge^{2}h_{*}\|$ by contradiction.
 Assume to the contrary that the inequality $R_{f}\geq(2m-1)(2m-2)\|\wedge^{2}h_{*}\|$ is strict somewhere.
@@
 here we take the metric $e^{-f}\cdot\langle\,$$\rangle$ on spinor bundle $S\otimes E$ and $C$ is a positive constant which is independent of $r$.
-Since $D_{f,E}\Psi=0$, combining equation (3.3) with equation (3.4), it follows that
+Since $D_{f,E}\Psi=0$, combining inequality (3.3) with inequality (3.4), it follows that
 $\begin{split}0\geq&\int_{M\times\mathbb{S}_{r}^{1}}\|\nabla\Psi\|^{2}e^{-f}+\frac{1}{4}\int_{M\times\mathbb{S}_{r}^{1}}\left(R_{f}-\sum_{i\neq j=1}^{2m-1}\tilde{\lambda}_{i}\tilde{\lambda}_{j}\right)\|\Psi\|^{2}e^{-f}\\ &-\frac{2m-1}{2r}\int_{M\times\mathbb{S}_{r}^{1}}\|{h}_{*}\|\|\Psi\|^{2}e^{-f}\\ \geq&\int_{M\times\mathbb{S}_{r}^{1}}\|\nabla\Psi\|^{2}e^{-f}+\frac{\delta_{0}}{4}\int_{U_{\delta}(x_{0})\times\mathbb{S}_{r}^{1}}\|\Psi\|^{2}e^{-f}\\ &-\frac{(2m-1)K}{2r}\int_{M\times\mathbb{S}_{r}^{1}}\|\Psi\|^{2}e^{-f}\\ \geq&\left(1-\frac{(2m-1)KC}{2r}\right)\int_{M\times\mathbb{S}_{r}^{1}}\|\nabla\Psi\|^{2}e^{-f}\\ &+\left(\frac{\delta_{0}}{4}-\frac{(2m-1)KC}{2r}\right)\int_{U_{\delta}(x_{0})\times\mathbb{S}_{r}^{1}}\|\Psi\|^{2}e^{-f}.\end{split}$
@@
 $(n-1)\left(k+1-\frac{n-6}{4}\right)=$ $2(n-1).$
-It follows from equation (3.7) taht
+It follows from equation (3.7) that
 $\int_{M}u^{\frac{n-2}{4}}e^{lf}|Au\nabla f+B\nabla u|^{2}=0,$
@@
 […]
-Plug $u\nabla f=-\frac{B}{A}\nabla u$ and equation (3.10) into equation (3.6) for k=0,m=0k=0,\,m=0, we obtain
+Plugging $u\nabla f=-\frac{B}{A}\nabla u$ and equation (3.10) into equation (3.6) for k=0,m=0k=0,\,m=0, we obtain
 $\left(2\frac{B}{A}-(n-1)\right)u\Delta u+\left(\frac{B^{2}}{A^{2}}-2\frac{B}{A}-\frac{1}{4}(n-1)(n-6)\right)|\nabla u|^{2}=0,$
@@
 ∎
-Remark 3.1.
-For $n\geq 2$, the Theorem 1.5 also holds if one replaces the operator norm $\|\wedge^{2}h_{*}\|$ by $\|h_{*}\|^{2}$.
+4. The Llarull type theorem in the form of $\mathbb{S}^{k}\times\mathbb{T}^{n-k}$
+In this section, we follow the approach of Hao-Shi-Sun [13] to establish the rigidity result of $\mathbb{S}^{k}\times\mathbb{T}^{n-k}$.
+We prove Theorem 1.11, specifically addressing the case $\mathbb{S}^{n-1}\times\mathbb{T}^{1}$.
+For the case $\mathbb{S}^{k}\times\mathbb{T}^{n-k}$, $2\leq n-k\leq n-3$, we use the torical symmetrization technique to achieve dimension reduction as described in [13, Remark 3.3](see also [27, Proposition 2.2]).
+Since Theorem 1.9 is crucial for the proof of Theorem 1.11, we first prove it.
+Proof of Theorem 1.9.
+By direct calculation, we have
+(4.1) $R(\bar{g})=R(g)+2\Delta f-|\nabla f|^{2}-\sum_{i=1}^{k}|\nabla f_{i}|^{2},$
+where $f=\sum_{i=1}^{k}f_{i}$.
+Since $R(\bar{g})\geq n(n-1)\|\wedge^{2}h_{*}\|$, $R(g)+2\Delta f-|\nabla f|^{2}\geq n(n-1)\|\wedge^{2}h_{*}\|$.
+By Theorem 1.5, we obtain there exists a constant $c>0$ such that $h:(M,c\cdot g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is an isometry and $f$ is a constant.
+Note $\|\wedge^{2}h_{*}\|=c$, by equation (4.1), $f_{i}$ is a constant for each $1\leq i\leq k$.
+∎
+Recall Theorem 1.11 for the specific case $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$.
+Theorem 4.1.
+Suppose $4\leq n\leq 7$ .
+Let $(M^{n},g)$ be a closed spin Riemannian manifold.
+If $h:(M,g)\to(\mathbb{S}^{n-1}\times\mathbb{S}^{1},g_{\mathbb{S}^{n-1}}+d\theta^{2})$ is a smooth map of non-zero degree such that $R(g)\geq(n-1)(n-2)\|\wedge^{2}h_{*}\|$, then there exists a constant $c>0$ such that $(M,g)$ is locally isometric to $(\mathbb{S}^{n-1}\times\mathbb{S}^{1},\frac{1}{c}\cdot(g_{\mathbb{S}^{n-1}}+d\theta^{2}))$.
+Proof.
+Since $h$ has non-zero degree, according to Sard’s theorem and the geometric measure theory, there exist a $\theta_{0}\in\mathbb{S}^{1}$ such that $0\neq[h^{-1}(\mathbb{S}^{n-1}\times\{\theta_{0}\})]\in H_{n-1}(M,\mathbb{Z})$ and a closed 2-sided area-minimizing hypersurface $\Sigma^{n-1}\in[h^{-1}(\mathbb{S}^{n-1}\times\{\theta_{0}\})]$, and $h_{1}\coloneqq P_{1}\circ h|_{\Sigma}:\Sigma\to\mathbb{S}^{n-1}$ has non-zero degree, where $P_{1}:\mathbb{S}^{n-1}\times\mathbb{S}^{1}\to\mathbb{S}^{n-1}$ is the canonical projection.
+By the second variation of stable minimal hypersurface, the associated Jacobi operator
+$J_{\Sigma}\coloneqq-\Delta_{\Sigma}-(\mathrm{Ric}_{g}(\nu,\nu)+|A_{\Sigma}|^{2})=-\Delta_{\Sigma}+\frac{1}{2}(R_{\Sigma}-R(g)-|A_{\Sigma}|^{2})$
+is non-negative, where $A_{\Sigma}$ is the second fundamental form of $\Sigma$ and $\nu$ is the unit normal vector of $\Sigma$. Let $\phi$ be the first eigenfunction such that
+(4.2) $\begin{split}J_{\Sigma}\phi=&-\Delta_{\Sigma}\phi-(\mathrm{Ric}_{g}(\nu,\nu)+|A_{\Sigma}|^{2})\phi\\ =&-\Delta_{\Sigma}\phi+\frac{1}{2}(R_{\Sigma}-R(g)-|A_{\Sigma}|^{2})\phi\\ =&\lambda\phi,\end{split}$
+where $\lambda\geq 0$ is the first eigenvalue.
+By the non-vanishing Theorem in [12, page 47], $\phi>0$.
+Since $\Sigma$ has trivial normal bundle, it is also a spin manifold.
+Now, we consider the manifold $(\Sigma\times\mathbb{S}^{1},g_{\phi}=g|_{\Sigma}+\phi^{2}d\theta^{2})$, which is spin, by direct calculation, we have on $\Sigma$
+$R(g_{\phi})=$ $R_{\Sigma}-\frac{2\Delta_{\Sigma}\phi}{\phi}$
+$=$ $R(g)+|A_{\Sigma}|^{2}+2\lambda$
+$\geq$ $(n-1)(n-2)\|\wedge^{2}(h|_{\Sigma})_{*}\|$
+$\geq$ $(n-1)(n-2)\|\wedge^{2}h_{1*}\|.$
+Since $h_{1}:\Sigma\to\mathbb{S}^{n-1}$ has non-zero degree, by Theorem 1.9, we deduce that $(\Sigma,g|_{\Sigma})$ is isometric to $(\mathbb{S}^{n-1},\frac{1}{c}\cdot g_{\mathbb{S}^{n-1}})$ for a constant $c>0$, $\|\wedge^{2}h_{1*}\|=c$ and $\phi$ is a constant.
+Utilizing equation (4.2), we obtain $\lambda=0$, $A_{\Sigma}=0$ and $\mathrm{Ric}_{g}(\nu,\nu)=0$.
+According to Proposition 3.2 in [4](see also Lemma 3.3 in [27]), we can construct a local foliation $\{\Sigma_{t}\}_{-\epsilon\leq t\leq\epsilon}$ in $M$ such that $\Sigma_{t}$ are of constant mean curvature and $\Sigma_{0}=\Sigma$, let $H_{t}$ denote the mean curvature of $\Sigma_{t}$.
+We claim that $H_{t}\equiv 0$ for any $t\in(-\epsilon,\epsilon)$.
+It suffices to prove this for nonnegative $t$, as the argument for the negative side is analogous.
+For contradiction, suppose that the assertion is not true, then there exists a $t_{0}>0$ such that $H_{t_{0}}=\epsilon_{0}>0$, since otherwise $H_{t}\leq 0$ and the area-minimizing property of $\Sigma$ will imply $H_{t}\equiv 0$.
+Let $\Omega_{t_{0}}$ denote the region enclosed by $\Sigma$ and $\Sigma_{t_{0}}$ i.e.,
+$\Omega_{t_{0}}:=\cup_{t\in[0,t_{0}]}\Sigma_{t}.$
+Consider the following functional
+(4.3) $E(\Omega)=\int_{\partial^{\ast}\Omega\setminus\Sigma}\mathrm{d}\mathcal{H}^{n-1}-\int_{\Omega_{t_{0}}}\epsilon_{1}(\chi_{\Omega}-\chi_{\bar{\Omega}})\mathrm{d}\mathcal{H}^{n},\,0<\epsilon_{1}<\epsilon_{0},$
+for $\Omega,\bar{\Omega}\in\mathcal{C}$ and $\Sigma\subset\bar{\Omega}$, where
+$\mathcal{C}=\{\Omega|\mbox{ all Caccioppoli sets }\Omega\subset\Omega_{t_{0}}\mbox{ and }\Omega\triangle\bar{\Omega}\Subset\overset{\circ}{\Omega}_{t_{0}}\}.$
+Due to $0<\epsilon_{1}<\epsilon_{0}$, we can find a Borel set $\hat{\Omega}$ minimizing the brane functional $E$ such that $\hat{\Sigma}=\partial\hat{\Omega}\setminus\Sigma$ is a smooth 2-sided hypersurface disjoint from $\Sigma$ and $\Sigma_{t_{0}}$.
+Let $\hat{\nu}$ and $\hat{A}$ denote the outwards pointing unit normal vector and the second fundamental form of $\hat{\Sigma}$, respectively.
+Since $\hat{\Sigma}$ is $E$-minimizing, it has constant mean curvature $\epsilon_{1}$ and it is $E$-stable.
+So the Jacobi operator
+$\hat{J}\coloneqq-\Delta_{\hat{g}}-(\mathrm{Ric}_{g}(\hat{\nu},\hat{\nu})+|\hat{A}|^{2})=-\Delta_{\hat{g}}+\frac{1}{2}(R_{\hat{\Sigma}}-R(g)-|\hat{A}|^{2}-\epsilon_{1}^{2})$
+is non-negative, where $\hat{g}$ is the induced metric of $\hat{\Sigma}$ from $(M,g)$.
+Taking the first eigenfunction $\hat{\phi}$ of $\hat{J}$ with respect to the first eigenvalue $\hat{\lambda}\geq 0$ and defining the metric $\tilde{g}=\hat{g}+\hat{\phi}^{2}d\theta^{2}$ on $\hat{\Sigma}\times\mathbb{S}^{1}$, then
+$R(\tilde{g})=R(g)+|\hat{A}|^{2}+2\hat{\lambda}+\epsilon_{1}^{2}\geq(n-1)(n-2)\|\wedge^{2}\hat{h}_{1*}\|+\epsilon_{1}^{2},$
+where $\hat{h}_{1}\coloneqq P_{1}\circ h|_{\hat{\Sigma}}:\hat{\Sigma}\to\mathbb{S}^{n-1}$, which has non-zero degree since $\hat{\Sigma}$ is homologous to $\Sigma$.
+By Theorem 1.9, we have $|\hat{A}|=0$ and $\epsilon_{1}=0$, which leads to a contradiction.
+Since $H_{t}$ vanishes for any $t\in(-\epsilon,\epsilon),$ there holds $\operatorname{Area}(\Sigma_{t})=\operatorname{Area}(\Sigma)$, which yields that $\Sigma_{t}$ is also area-minimizing.
+Replacing $\Sigma$ by $\Sigma_{t}$ and combining with Theorem 1.9, $\Sigma_{t}$ is totally geodesic and has vanished normal Ricci curvature.
+In the following, we establish the existence of a local isometry $\varPhi:\Sigma\times\mathbb{R}\to M$.
+Let $\tilde{\nu_{t}}$ and $\tilde{V_{t}}$ denote the outwards pointing unit normal vector field and normal variation vector field of the foliation $\{\Sigma_{t}\}_{-\epsilon\leq t\leq\epsilon}$, respectively.
+Assume $\varPhi:\Sigma\times(-\epsilon,\epsilon)\to M$ is the flow generated by $\tilde{V_{t}}$.
+It follows that $\varPhi$ is an embedding in a small neighborhood of $\Sigma$ and the pull-back of the metric $g$ is
+$\varPhi^{\ast}(g)=\psi^{2}dt^{2}+\varPhi^{\ast}_{t}(g_{t}),$
+where $\psi=\langle\tilde{V_{t}},\tilde{\nu_{t}}\rangle>0$ is the lapse function (see definition in (3.3) in [4]) and $g_{t}$ is the induced metric on $\Sigma_{t}$ from $g$.
+Since $\Sigma_{t}$ is totally geodesic, we have $\partial_{t}\varPhi^{\ast}_{t}(g_{t})=2\psi A_{\Sigma_{t}}=0$, which implies that $\varPhi^{\ast}_{t}(g_{t})=g|_{\Sigma}$.
+Additionally, the stability condition yields $\Delta_{\Sigma_{t}}\psi-|A_{\Sigma_{t}}|^{2}\psi-\operatorname{Ric}(\nu_{t},\nu_{t})\psi=-\Delta_{\Sigma_{t}}\psi=0,$ then $\psi(\cdot,t)\equiv\mathrm{constant}$.
+Let
+$r(\cdot,t)=\int_{0}^{t}\psi(\cdot,s)ds,$
+then
+$\varPhi^{\ast}(g)=dr^{2}+g|_{\Sigma}.$
+Hence, $\varPhi:\Sigma\times(-\epsilon,\epsilon)\to M$ is a local isometry.
+Through a continuous argument as Proposition 3.8 in [4], we conclude that there exists a local isometry $\varPhi:\Sigma\times\mathbb{R}\to M$.
+Since $(\Sigma,g|_{\Sigma})$ is isometric to $(\mathbb{S}^{n-1},\frac{1}{c}\cdot g_{\mathbb{S}^{n-1}})$ for a constant $c>0$, by scaling the fact $\mathbb{R}$, $(M,g)$ is isometrically covered by $(\mathbb{S}^{n-1}\times\mathbb{R},\frac{1}{c}\cdot(g_{\mathbb{S}^{n-1}}+d\theta^{2}))$.
+∎